# Comparison of the Migdal transition probabilities in electron-atom inelastic cross sections

### v3

@@ Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability in the Born approximation. The resultant errors are around $20\ \%$ for Ar and $60\ \%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated by the outermost shell electrons. The error for Xe atom will become better for high energy because of the approximation and lack of data.
@@ 1 Introduction
The Migdal effect migdal1939ionizatsiya ; migdal1941ionizacia ; Landau:1991wop is effective for searching small energy signals from dark matter. The Migdal effect is a inelastic process of dark matter-nucleus scattering and causes the excitation/ionization of electrons from the nuclear recoil. The excitation/ionization of electrons are electromagnetic signals that are easier to detect than heat in the case of elastic scattering. The Migdal effect is not observed in nuclear scattering, but, observed in $\alpha$ decay Rapaport:1975zza ; Rapaport:1975zz ; fischbeck1975spectroscopy ; fischbeck1977angular and $\beta$ decay boehm1954internal ; berlovich1965investigation ; couratin2012first ; Lienard:2015joa ; Fabian:2018pin . Now, experiments for detecting the Migdal effect from nuclear recoils are proceeding Nakamura:2020kex ; Araujo:2022wjh . Recently, the Migdal effect is studied or used in many papers baur1983ionisation ; vegh1983multiple ; wauters1997recoil ; berakdar2001electronic ; Vergados:2005dpd ; Moustakidis:2005gx ; Ejiri:2005aj ; talman2006excitations ; Bernabei:2007jz ; liertzer2012multielectron ; Vergados:2013raa ; pindzola2014neutron ; Sharma:2017fmo ; Ibe:2017yqa ; Dolan:2017xbu ; LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr ; Coskuner:2018are ; EDELWEISS:2019vjv ; Tvrznikova:2019tgx ; CDEX:2019hzn ; XENON:2019zpr ; Baxter:2019pnz ; Essig:2019xkx ; Liang:2019nnx ; Kozaczuk:2020uzb ; SENSEI:2020dpa ; Liu:2020pat ; Nakamura:2020kex ; Pindzola:2020whx ; Kahn:2020fef ; Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu ; LUX:2020yym ; Liang:2020ryg ; He:2020sat ; Flambaum:2020xxo ; Akerib:2021pfd ; Collar:2021fcl ; Bell:2021zkr ; Liao:2021yog ; Menshikov:2021jxa ; Acevedo:2021kly ; COSINE-100:2021poy ; Berlin:2021zbv ; CDEX:2021cll ; Wang:2021oha ; Bell:2021ihi ; Abbamonte:2022rfh ; Chatterjee:2022gbo ; EDELWEISS:2022ktt ; Angevaare:2022zhy ; Campbell-Deem:2022fqm ; Araujo:2022wjh ; DarkSide:2022dhx ; Blanco:2022pkt ; Cox:2022ekg ; GlobalArgonDarkMatter:2022ppc ; Tomar:2022ofh ; Adams:2022zvg ; Berghaus:2022pbu ; Li:2022acp ; NEWS-G:2023qwh ; DarkSide-50:2023fcw ; SuperCDMS:2023sql ; Bell:2023uvf ; Tilly:2023fhw ; AtzoriCorona:2023ais ; Qiao:2023pbw ; Xu:2023wev ; LZ:2023poo ; PandaX:2023xgl ; Yun:2023huf ; Gu:2023pfg ; Li:2023xkf ; Herrera:2023xun ; Kim:2023lkb ; Dzuba:2023zcq ; SENSEI:2023zdf ; Liang:2024tef ; He:2024hkr ; Balan:2024cmq ; MIGDAL:2024alc ; Kang:2024kec ; Ibarra:2024mpq ; Lee:2024wzd , which includes that the comparison LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr to bremsstrahlung of photon Kouvaris:2016afs ; Millar:2018hkv , theoretical comparison to the dark matter-electron scattering Baxter:2019pnz ; Essig:2019xkx , the Migdal-photoabsorption relation Liu:2020pat , formulation for semiconductor response Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu , and the semi-inclusive calculation Cox:2022ekg .
@@
The energy eigenstates of atom are given in Sec. 2. The formalism uses the center of mass coordinates and is useful for treating long-range interactions. The interaction potential is derived in Sec. 3. The interactions are approximately described by the electron-electron interaction. Therefore, the shape of potential is the Coulomb one. The total excitation and ionization cross section is calculated in Sec. 4. Especially, the cross section is expressed by the Migdal transition factor of dark matter-atom scattering for single electron excitation/ionization with the dipole approximation given in Ref. Ibe:2017yqa . Then, the result is compared with a database value which is a combination of experimental and theoretical data. The natural unit is used in this paper.
@@ 3 Interaction potential
The interaction with the nucleus can be ignored approximately when the excitation and ionization processes are considered. This is because the potential from the nucleus interactions depend on the coordinates as $\vec{r}-\vec{r}_{n}$ and $\vec{r}_{n}$ is related to the relative coordinates $\vec{r}_{i}$ with the suppression factor of $\approx m_{e}/m_{A}$.
@@ 4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in}\rangle_{I/F}$. The incoming electron states can be taken as plane waves because the atom is regarded to be neutral far from the atom even the atom is ionized. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
@@
×$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I}\rangle$
×$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F}|\vec{r}\rangle\phi_{E_{e},I}(\vec{r})e^{i\vec{q}_{u}\cdot\vec{r}}.$ (7)
Here, $\vec{p}_{{\rm CM},u,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ because of the $u$-channel process. The integrations by $\vec{r}_{\rm in}$ and $\vec{r}$ are actually the Fourier transformations of the projected single-electron wave functions. Assuming that the plane wave $\langle\vec{r}|\vec{p}\rangle$ is approximately equal to the Coulomb wave $\phi(\vec{r})$ (see Ref. gozem2015photoelectron for the discussion of the differences), the above equation becomes
@@
In the cross section, the possible final states $F$ and spins are summed or averaged. The electron spins of final states $s_{i}^{F}$ are included in the label $F$. Note that the sum includes the integration of ionized electron energy. This equation is reduced as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{2\pi\alpha^{2}m_{e}^{2}}{\vec{p}_{{\rm in},I}\,^{2}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{q^{2}_{\rm max}}{q^{2}_{\rm min}}+\left|z_{{\rm in}I}\right|^{2}\log\frac{2m_{e}E_{{\rm kin},I}-q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}-q_{\rm max}^{2}}\right.$
@@
$\sigma(E_{{\rm kin},I})=$ $\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{E_{{\rm kin},I}-\Delta E}{\Delta E}\right.$
$\left.-{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{1-\frac{\Delta E}{E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{\Delta E}{E_{{\rm kin},I}}}\right)\right].$ (14)
To obtain Eq. (13), the average/sum of magnetic quantum number for the initial/final states are taken, which makes the factor $\left|z_{FI}(\theta_{qr})\right|^{2}$ isotropic for the angular direction $\left|z_{FI}(\theta_{qr})\right|^{2}\to\left|z_{FI}\right|^{2}$ Ibe:2017yqa . Also, the fact that the transition matrix element does not change the spins is used. In Eq. (14), the relation $\sum_{F}\left|z_{FI}\right|^{2}=\sum_{F}\left|z_{{\rm in}I}\right|^{2}$ is used. The final states of $z_{FI}z_{{\rm in}I}^{*}$ are related by the energy conservation. The first term is actually the cross section by the Coulomb potential without the identification factor of the interacting electrons $1/2!$.
The averaged value of transition probabilities $\sum|z_{FI}|^{2}$ is actually the transition probabilities for the Migdal effect given in Ref. Ibe:2017yqa which uses the Flexible Atomic Code (FAC, cFAC) doi:10.1139/p07-197 . cFAC calculates the probabilities with the Dirac-Hartree-Fock method and a universal central potential.
@@
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the calculation error of ionization is around $20\ \%$ from the database value in 300 eV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around $60\ \%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the error of ionization is around 60 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This error dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term decreases the cross section for low energy and negligible for high energy. Neglected the interference term, the overestimation near the threshold energy is consistent with Ref. ariyasinghe2004total which compares to the Bethe-Born theory inokuti1971inelastic . Note that the probability is dominated by the excitation/ionization from the outermost shell, i.e., $3p$ states for Ar and $5p$ states for Xe.

### v5

@@ Abstract
The Migdal transition probabilities for dark matter scattering are compared to the total single-electron inelastic cross sections of electron-atom scattering for isolated Ar and Xe. The comparison is done by expressing the electron-atom scattering cross section by connecting the Migdal probability. The resultant differences are around $30\ \%$ for Ar and $80\ \%$ for Xe in $\sim$ 1 keV of incoming electron energy. The transition is dominated from the $3p$ shell electrons for Ar. For Xe, $5p$ states dominate the contribution, but, $4d$ states give $\sim 40\ \%$ contribution at 1 keV.
@@ 1 Introduction
The Migdal effect migdal1939ionizatsiya ; migdal1941ionizacia ; Landau:1991wop is effective for searching small energy signals from dark matter. The Migdal effect is a inelastic process of dark matter-nucleus scattering and causes the excitation/ionization of electrons from the nuclear recoil. The excitation/ionization of electrons returns electromagnetic signals that are easier to detect than heat in the case of elastic scattering. The Migdal effect is not observed in nuclear scattering, but, observed in $\alpha$ decay Rapaport:1975zza ; Rapaport:1975zz ; fischbeck1975spectroscopy ; fischbeck1977angular and $\beta$ decay boehm1954internal ; berlovich1965investigation ; couratin2012first ; Lienard:2015joa ; Fabian:2018pin . Now, experiments for detecting the Migdal effect from nuclear recoils are proceeding Nakamura:2020kex ; Araujo:2022wjh . Recently, the Migdal effect is studied or used in many papers baur1983ionisation ; vegh1983multiple ; wauters1997recoil ; berakdar2001electronic ; Vergados:2005dpd ; Moustakidis:2005gx ; Ejiri:2005aj ; talman2006excitations ; Bernabei:2007jz ; liertzer2012multielectron ; Vergados:2013raa ; pindzola2014neutron ; Sharma:2017fmo ; Ibe:2017yqa ; Dolan:2017xbu ; LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr ; Coskuner:2018are ; EDELWEISS:2019vjv ; Tvrznikova:2019tgx ; CDEX:2019hzn ; XENON:2019zpr ; Baxter:2019pnz ; Essig:2019xkx ; Liang:2019nnx ; Kozaczuk:2020uzb ; SENSEI:2020dpa ; Liu:2020pat ; Nakamura:2020kex ; Pindzola:2020whx ; Kahn:2020fef ; Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu ; LUX:2020yym ; Liang:2020ryg ; He:2020sat ; Flambaum:2020xxo ; Akerib:2021pfd ; Collar:2021fcl ; Bell:2021zkr ; Liao:2021yog ; Menshikov:2021jxa ; Acevedo:2021kly ; COSINE-100:2021poy ; Berlin:2021zbv ; CDEX:2021cll ; Wang:2021oha ; Bell:2021ihi ; Abbamonte:2022rfh ; Chatterjee:2022gbo ; EDELWEISS:2022ktt ; Angevaare:2022zhy ; Campbell-Deem:2022fqm ; Araujo:2022wjh ; DarkSide:2022dhx ; Blanco:2022pkt ; Cox:2022ekg ; GlobalArgonDarkMatter:2022ppc ; Tomar:2022ofh ; Adams:2022zvg ; Berghaus:2022pbu ; Li:2022acp ; NEWS-G:2023qwh ; DarkSide-50:2023fcw ; SuperCDMS:2023sql ; Bell:2023uvf ; Tilly:2023fhw ; AtzoriCorona:2023ais ; Qiao:2023pbw ; Xu:2023wev ; LZ:2023poo ; PandaX:2023xgl ; Yun:2023huf ; Gu:2023pfg ; Li:2023xkf ; Herrera:2023xun ; Kim:2023lkb ; Dzuba:2023zcq ; SENSEI:2023zdf ; Liang:2024tef ; He:2024hkr ; Balan:2024cmq ; MIGDAL:2024alc ; Kang:2024kec ; Ibarra:2024mpq ; Lee:2024wzd , which includes that the comparison LUX:2018akb ; Bell:2019egg ; GrillidiCortona:2020owp ; vonKrosigk:2020udi ; XMASS:2022tkr to bremsstrahlung of photon Kouvaris:2016afs ; Millar:2018hkv , theoretical comparison to the dark matter-electron scattering Baxter:2019pnz ; Essig:2019xkx , the Migdal-photoabsorption relation Liu:2020pat , formulation for semiconductor response Knapen:2020aky ; Knapen:2021bwg ; Liang:2022xbu , and the semi-inclusive calculation Cox:2022ekg .
@@
The energy eigenstates of atom are given in Sec. 2. The formalism uses the center of mass coordinates and is useful for treating long-range interactions. The interaction potential is derived in Sec. 3. The interactions are approximately described by the electron-electron interaction. Therefore, the shape of potential is the Coulomb one. The total excitation and ionization cross section is calculated in Sec. 4. Especially, the cross section is expressed by the Migdal transition factor of dark matter-atom scattering for single electron excitation/ionization with the dipole approximation given in Ref. Ibe:2017yqa . Then, the result is compared with a database value which is a combination of experimental and theoretical data and other theoretical calculations. The natural unit is used in this paper.
@@ 3 Interaction potential
The interaction with the nucleus can be ignored approximately when the excitation and ionization processes are considered. This is because the potential from the nucleus interactions depend on the coordinates as $\vec{r}-\vec{r}_{n}$ and $\vec{r}_{n}$ is related to the relative coordinates $\vec{r}_{i}$ with the suppression factor of $\approx m_{e}/m_{n}$.
@@ 4 Electron-atom inelastic scattering cross section
The initial state $|\Phi_{I}\rangle$ and final state $|\Phi_{F}\rangle$ for asymptotic states are given by $|\Phi_{I/F}\rangle=\sqrt{2m_{A}}\sqrt{2m_{e}}|\Psi_{A}\rangle_{I/F}|\vec{p}_{\rm in},s_{\rm in}\rangle_{I/F}$. The incoming electron states can be taken as plane waves because the atom is regarded to be neutral far from the atom. Then, the $T$-matrix by a $t$-channel potential $V_{t}(\vec{r}_{i}+\vec{r}_{\rm CM}-\vec{r}_{\rm in})$ becomes
@@
×$\times\int d^{3}\vec{r}_{\rm in}\phi_{E_{e},F}^{\dagger}(\vec{r}_{\rm in})e^{-i\vec{q}_{u}\cdot\vec{r}_{\rm in}}\langle\vec{r}_{\rm in}|\vec{p}_{{\rm in},I},s_{{\rm in},I}\rangle$
×$\times\int d^{3}\vec{r}\langle\vec{p}_{{\rm in},F},s_{{\rm in},F}|\vec{r}\rangle e^{i\vec{q}_{u}\cdot\vec{r}}\phi_{E_{e},I}(\vec{r}).$ (7)
Here, $\vec{p}_{{\rm CM},u,F}=\vec{p}_{{\rm CM},F}-\vec{p}_{i,F}+\vec{p}_{{\rm in},F}$ because of the $u$-channel process. The integrations by $\vec{r}_{\rm in}$ and $\vec{r}$ are actually the Fourier transformations of the projected single-electron wave functions. Assuming that the plane wave $\langle\vec{r}\,|\vec{p},s\rangle$ is approximately equal to the Coulomb wave $\phi(\vec{r})$ (see Ref. gozem2015photoelectron for the discussion of the differences), the above equation becomes
@@
In the cross section, the possible final states $F$ and spins are summed or averaged. The electron spins of final states $s_{i}^{F}$ are included in the label $F$. The sum of the initial spin for atomic electrons are included in $Z_{FI}$ Ibe:2017yqa . Note that the sum includes the integration of ionized electron energy. This equation is reduced as
$\sigma(\vec{p}_{{\rm in},I}\,^{2})=$ $\frac{4\pi\alpha^{2}m_{e}^{2}}{\vec{p}_{{\rm in},I}\,^{2}}\sum_{F}\left[\left|z_{FI}\right|^{2}\log\frac{q^{2}_{\rm max}}{q^{2}_{\rm min}}+\left|z_{{\rm in}I}\right|^{2}\log\frac{2m_{e}E_{{\rm kin},I}-q_{\rm min}^{2}}{2m_{e}E_{{\rm kin},I}-q_{\rm max}^{2}}\right.$
@@
$\sigma(E_{{\rm kin},I})=$ $\frac{2\pi\alpha^{2}m_{e}}{E_{{\rm kin},I}}\sum_{F}\left[\left(\left|z_{FI}\right|^{2}+\left|z_{{\rm in}I}\right|^{2}\right)\log\frac{E_{{\rm kin},I}-\Delta E}{\Delta E}\right.$
$\left.-2{\rm Re}\left[z_{FI}z_{{\rm in}I}^{*}\right]\left(\sin^{-1}\sqrt{1-\frac{\Delta E}{E_{{\rm kin},I}}}-\sin^{-1}\sqrt{\frac{\Delta E}{E_{{\rm kin},I}}}\right)\right].$ (14)
To obtain Eq. (13), the average/sum of magnetic quantum number for the initial/final states are taken, which makes the factor $\left|z_{FI}(\theta_{qr})\right|^{2}$ isotropic for the angular direction $\left|z_{FI}(\theta_{qr})\right|^{2}\to\left|z_{FI}\right|^{2}$ Ibe:2017yqa . Also, the fact that the transition matrix element does not change the spins is used. The final states of $z_{FI}z_{{\rm in}I}^{*}$ are related by the energy conservation as $E_{{\rm in},F}\approx E_{{\rm kin},I}-\Delta E$. The log term shows the kinematical constraint from the phase space of $E_{e,F}$ as $E_{{\rm kin},I}-2\Delta E>0$. The $\left|z_{FI}\right|^{2}$ part in the first term is actually the scattering cross section by the Coulomb potential without the identification factor of the interacting electrons $1/2!$.
The averaged value of $|z_{FI}|^{2}$ is actually the transition probabilities for the Migdal effect given in Ref. Ibe:2017yqa which uses the Flexible Atomic Code (FAC, cFAC) doi:10.1139/p07-197 for the calculation. cFAC calculates the probabilities with the Dirac-Hartree-Fock method and a universal central potential.
@@
The calculation result of $\sigma(E_{{\rm kin},I})$ is shown in Fig. 1. For Ar, the difference of ionization cross section is around $30\ \%$ from the database value in 5 keV $\lesssim E_{{\rm kin},I}<10$ keV. That of excitation has local maximum of around $60\ \%$ at 100 eV and gradually decreasing by increasing $E_{{\rm kin},I}$. For Xe, the difference of ionization cross section is around 80 % at 1 keV and that of excitation is roughly factor 3 at 1 keV. This gap dependency is consistent with that the Born approximation is a good approximation in the keV energy region bonham1974high ; garcia1999electron . The interference term is effective at low energy and negligible at high energy. Note that the probability is almost dominated by the excitation/ionization from the outermost shell. For Ar, $3p$ states dominate the contribution and $2p$ and $3s$ states give $\sim 1\ \%$ contribution at $E_{{\rm kin},I}=10$ keV. For Xe, $5p$ states dominate up to $\sim$ 100 eV, but, $4d$ states have $\sim 40\ \%$ contribution at $E_{{\rm kin},I}=1$ keV.
5 Comparison with other results
The inelastic cross section is also calculated by using the Bethe theory inokuti1971inelastic ; inokuti1975total ; inokuti1981oscillator ; salvat2022inelastic . The obtained formula of inelastic cross section for Ar is inokuti1975total
$\sigma_{\rm Inokuti}(E_{{\rm kin},I})=4\pi a_{0}^{2}\frac{\rm Ryd}{E_{{\rm kin},I}}\left[M_{\rm Inokuti}^{2}\log\left(4\frac{E_{{\rm kin},I}}{\rm Ryd}\right)+M_{\rm Inokuti}^{2}\log c_{\rm Inokuti}+\gamma_{\rm Inokuti}\frac{\rm Ryd}{E_{{\rm kin},I}}\right],$ (1)
where $a_{0}=1/(m_{e}\alpha)$ is the Bohr radius, Ryd = 13.605 eV is the Rydberg energy, $\gamma_{\rm Inokuti}=Z(-7/4+\log(B/E_{{\rm kin},I}))$ for electron, and $B$ is the average binding energy which is taken as 13 eV because the 3p states dominate the signal. $M_{\rm Inokuti}^{2}=5.4755$ and $M_{\rm Inokuti}^{2}\log c_{\rm Inokuti}=-1.063$ are taken from the Hartree-Fock calculation (see Ref. inokuti1975total ). Ref. inokuti1975total ; inokuti1981oscillator calculated the ionization cross section from hydrogen ($Z=1$) to strontium ($Z=38$).
For Xe, the result of Ref. salvat2022inelastic is used. It shows
$\sigma_{\rm Salvat}(\beta^{2})=\frac{2\pi\alpha^{2}}{m_{e}\beta^{2}}\left[M_{\rm Salvat}^{2}\left(\log\frac{\beta^{2}}{1-\beta^{2}}-\beta^{2}\right)+C_{\rm Salvat}\right],$ (2)
where $\beta$ is the velocity of incoming electron and the values of $M_{\rm Salvat}^{2}$ and $C_{\rm Salvat}$ are taken from the figure in Ref. salvat2022inelastic . This calculation used the shell corrections, self-consistent Dirac-Hartree-Fock-Slater potential, independent-electron approximation, and relativistic plane-wave Born approximation. They reported less than 1 % precision for $E\gtrsim 100$ MeV and $\sim 8$ precision for $E\sim 5$ MeV for high $Z$ atoms and good agreement with the results of Ref. inokuti1975total ; inokuti1981oscillator .
Figure 2: The comparison of total inelastic cross section $\sigma(E_{{\rm kin},I})$. The calculation results using Ref. Ibe:2017yqa are shown by the solid lines and the comparison values are shown by the dashed and dotted lines. (top:) For Ar case. The dotted lines are from the theoretical calculation in Ref. inokuti1975total with the Hartree-Fock calculation. The database value is taken from Phelps Phelps:2024 . (bottom:) For Xe case. The dotted lines are from the theoretical calculation in Ref. salvat2022inelastic and not convergent at low energy region because of its formula. The database value is taken from Morgan Morgan:2024 .
The comparison result is shown in Fig. 2. For both comparisons, this result shows smaller value than the database value and Ref. inokuti1975total ; salvat2022inelastic . Also, the result from Ref. salvat2022inelastic gets consistent parameters with that from Ref. inokuti1975total . The line of Ref. salvat2022inelastic aims over MeV region with heavy incoming particles and does not converge at low energy region because of its shape of formula. Their result will be applied here because electrons are not fast in keV.
The binding energies and wave functions of electrons are calculated with electron correlations, e.g., Ref. Baxter:2019pnz shows $\sim 5$ precision for binding energies which is better than Ref. Ibe:2017yqa . Also, Ref. Cox:2022ekg shows $\sim 20$ differences of ionization probabilities for xenon 5s and argon 3p states with a 20 keV ionized electron. These uncertainties are significant for outer shell states and this comparison method actually relies on the precision of outermost shell states. Note that the uncertainty becomes much smaller for inner shell electrons. Density functional theory may provide better precision of binding energies, but, it does not provide physical wave functions.
In addition, this scattering formulation approximates the electron correlations, but this uncertainty is difficult to estimate. The use of $u$-channel potential will be valid for $E_{{\rm kin},I}\gtrsim 1$ keV electron because faster electrons feel less effect of atomic electron clouds than slower ones.